IF three persons namely A,B and C can do a work in 12 days,18 days and 19days respectively . If they all do the work for 10 days and then another person D completed the remaining work in 14 days , In how many days D alone can finish That work ?

Grade:10

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Rate of doing work of A, B, C
w/12, w/18, w/19

Work in 10 days = w ( 1/12 + 1/18+ 1/19)= w*0.2

Time req by D to finish 0.8 w = 14 days

Time taken to complete w work by D = 14/0.8 = 17.5 days

Regards